OpenStudy (anonymous):

-3.4-4x-3.6-x=-4x+0.1

OpenStudy (anonymous):

-7.1

OpenStudy (anonymous):

\[-3.4 - 4x-3.6-x=-4x+0.1\]First combine like terms:\[-7 - 5x=-4x+0.1\]Add 5x to both sides:\[-7=x+0.1\]Subtract 0.1 from both sides:\[-7.1=x\]